TEXcad Version 2.8

Copyright \begin{matrix}c \\ ◯ \end{matrix} Georg Horn

October 1990

Contents

1 Introduction 2

2 Using TEXcad 3 2.1 Contents of the disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Starting TEXcad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 The menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.5 The message window, mouse and key functions . . . . . . . . . . . . . . . . . . . . . . . . 5 2.6 The drawing window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 Drawing pictures 6 3.1 Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Boxes, dashed and filled boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Lines and arrows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.3.1 Ordinary lines and arrows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3.2 EMLines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.4 Circles and filled circles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.5 Ellipses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.6 Bezier-Curves . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Editing pictures 8 4.1 Picking objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Editing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3 Copying and moving objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4 Deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.5 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Utilities 10

5.1 Redraw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5.2 The snap function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5.3 Zoom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5.4 Moving the window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6 Options 11

6.1 Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6.2 Snap, Zoom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6.3 Dimensional units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6.4 EMLines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6.5 Loading and saving options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

7 Loading and saving pictures 12

7.1 Picture file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

8 Ending the run 13

A The file gh.bat 13

B The file gh.tex 14

C Licensing conditions 14

1 Introduction

TEXcad is a drawing program for LaTEX which allows the objects available in the LaTEX picture environment to be drawn and edited, its output is a sequence of LaTEX picture commands which can be inserted into a LaTEX document to generate the drawing. These picture commands are best brought into the document by means of the \input command so that, if necessary, the files can be read in to TEXcad and edited again.

TEXcad was originally implemented as a student project at the University of Coblenz under the supervision of Dr W. Schiffmann. Since then it gone through several improved versions and is included in Eberhard Mattes’ emTEX package (see also appendix C). The list of the preceding versions that appeared at this place is now in changes.doc:

So eine Arbeit wird eigentlich nie fertig, man muß sie für fertig erklären, wenn man nach Zeit und Umständen das möglichste getan hat. (J. W. Goethe, Italienische Reise, 16.3.1787.)

Like all programs, TEXcad is sure to contain bugs and could be improved: error reports and suggestions for improvement are welcome.

2 Using TEXcad

TEXcad is written in Turbo Pascal V5 and runs on all IBM PC/XT/AT machines and compatibles with a graphics adapter for which there is a Turbo Pascal driver. The Borland drivers are supplied with the program package. For efficient working, your computer should have a mouse, the mouse driver must then be loaded before starting TEXcad.

2.1 Contents of the disk

The disk has the following files:

texcad.exe the TEXcad program
texcad.opt this file contains the parameters described in sections 2.3 and 6

gh.bat a batch file incorporating the DOS commands to run the editor, LaTEX, Preview

and TEXcad (see appendix A)

ask.exe called from gh.bat to get input from the user
gh.tex a TEX file for testing picture commands generated by TEXcad or other LaTEX files
*.bgi the Borland Graphics Interface drivers

emlines.sty a style file which allows lines to be drawn at any angle when the emTEX drivers are

used for output: these lines will be called EMLines in what follows.

emlines2.sty a new version of emlines.sty to be used with the emTEX drivers V1.3j and later

(See section 3.3.2).

tcman.dvi this manual.
texcad.ver list of preceeding versions.

2.2 Installation

Copy the files texcad.exe, texcad.opt, gh.bat, ask.exe and the *.bgi files into the same directory as LaTEX and Preview, for instance, \emtex. If you already have Turbo Pascal on your disk then the *.bgi files need not be copied, you must just tell TEXcad the directory containing them (see sections 2.3 and 6). Copy emlines*.sty and gh.tex into the same directory as the other .sty files, for example,

\emtex\texinput.

When delivered with emTEX, the files are distributed in a .zip file and must be unpacked with pkunzip before they can be used. In this case, you must already have done this using the instructions provided with emTEX as otherwise you would not be reading this manual.

2.3 Starting TEXcad

TEXcad is started with the command texcad <filename>, the optional argument filename is a file containing a LaTEX picture environment – this picture will be loaded immediately (see also section 7). But first, the mouse driver will be initialized, if no mouse is present, TEXcad issues an apropriate message. TEXcad then looks in the current directory and then in the directory containing texcad.exe for the parameter file texcad.opt. This file contains the following arguments – if the file is not present, defaults are taken:

Parameter  Default
Directory for loading and storing picture files none
Extension for picture files  .pic
Directory for loading and storing macro files none
Extension for macro files  .mac
Directory for loading graphic drivers  none
Snap function  off
Magnification  1
Line thickness for EMLines  0.4pt
Draw EMLines  off

                    \unitlength setting  1mm

The parameters can be changed with the menu Options and the new values saved in texcad.opt with the Save entry in this menu (see section 6).

Finally TEXcad loads the Borland graphics driver: these are sought first in the directory given in texcad.opt, if no directory is given there, then the current directory is searched. If the driver cannot be found then you are asked to enter the name of the directory containing them or, by pressing ENTER, to leave the program.

You are recommended to put each LaTEX document you create in a separate subdirectory, set the parameters as required and save them in this subdirectory. If the parameters are unlikely to be changed for other documents, you can copy the option file to the directory containing texcad.exe. It is also possible to create a parameter file for each picture, TEXcad will search for a parameter file path\name.opt, when loading a picture path\name.pic

After a successful start, the screen will be divided as follows:

Drawing window Message window

Menu window

The message window shows the version number and a copyright notice, and the menu window displays the main menu:

                                  MAIN-MENU

Draw Edit Redraw Snap Zoom Window Options Save Load Quit

2.4 The menus

The highlight bar in the menu can be moved either with the mouse or the cursor keys and the selection made by pressing the left mouse button or the ENTER key. Some of the menu items will then display a submenu for you to make a further choice. The title of the menu will always appear at the top of the window. You can leave a submenu by pressing the right mouse button or the ESC key – except that the main menu can be left (and the program ended) only by selecting Quit from it.

2.5 The message window, mouse and key functions

This window is used to display messages from TEXcad and for input from the user. There are four kinds of messages:

1. A request for text input: a file name, for instance. During text entry the following keys and mouse buttons may be used:

Left and right arrow keys move the cursor one character to the left or right

Insert toggles between insert and overwrite mode

Delete removes the character under the cursor

Backspace removes the character to the left of the cursor

ESC or right button ends input and terminates the operation

ENTER or left button ends input and returns the entered text to the program

2. A request for a point to be marked in the drawing window. To do this the cross hairs must be moved with the mouse or the arrow keys^{1} to the desired position and the left mouse button or ENTER pressed. If an error is made, the operation can be terminated by pressing the right mouse button.

While marking points, the display window can be moved upwards, downwards, to the left or to the right by pressing the keys E, X, S and D. The snap function can be toggled on and off with the P key.

3. Error messages – to continue, press ENTER or the left mouse button.

4. Yes/No queries which must be answered by pressing Y, ENTER or the left button for “Yes”, or N, ESC or the right button for “No”.

1Ctrl-arrow keys move the cursor in greater steps

2.6 The drawing window

The top and righthand edges of the window carry a ruler against which the horizontal and vertical position of the cross hairs can be read. The distance between two divisions is the value set in menu

Options/Unitlength.

3 Drawing pictures

Selecting Draw from the main menu presents you with a menu of the objects which can be drawn:

                                    DRAW

Text Framebox Dashbox Filled Box Line Vector Circle Filled Circle Oval Half/Quarter Oval Bezier

3.1 Text

The first item on the Draw menu allows you to enter text into the picture. After you have selected the reference point for the text, the upper part of the window will be cleared and you can enter any LaTEX text you wish. When you have done this, you will have to select the text’s horizontal and vertical positioning with respect to the reference point. The choices for horizontal positioning are:

Left: the lefthand end of the text will be at the reference point, that is, the text lies to the right of the reference point.

Right: the righthand end of the text will be at the reference point, that is, the text lies to the left of the reference point.

Center: the text will be centred horizontally.

The choices for vertical positioning are:

Top: the top of the text box will be at the reference point, that is, the text lies below the reference point.

Bottom: the bottom of the text box will be at the reference point, that is, the text lies above the reference point.

Center: the text will be centred vertically.

When these choices have been made, the normal picture will reappear with the position of the text shown by a T and you will return to the Draw menu.

3.2 Boxes, dashed and filled boxes

Selecting Framebox from the menu draws a box with solid lines, Dashbox gives a box with broken lines and Filled Box gives a filled in box.

To draw a box you must first select the bottom lefthand corner of the box and the righthand top corner. If you have selected Dashbox then you must also give the length of the dashes by marking two points at the beginning and end of the desired length. Having done this you can now enter some text to go in the box – if you do not want text in the box then you must enter an empty string. Text is entered as described in section 3.1 except that the positioning of the text in the box is as follows:

Left: the lefthand end of the text will start at the left edge of the box, that is, the text is left justified in the box.

Right: the righthand end of the text will be against the right edge of the box, that is, the text is right justified in the box.

Center: the text will be centred horizontally in the box.

For vertical positioning

Top: the top of the text box will be against the top edge of the box, that is, the text is at the top of the box.

Bottom: the bottom of the text box will be against the bottom of the box, that is, the text is at the bottom of the box.

Center: the text will be centred vertically in the box.

A box with text in it will have the letters BT in it.

3.3 Lines and arrows

The menu item Line draws lines and the item Vector draws lines with arrow heads. The line is drawn by marking the starting point first and then the end point. You can then draw further lines by marking further endpoints, the previous endpoint is always taken as the starting point. The drawing of lines is aborted with ESC or right button.

3.3.1 Ordinary lines and arrows

As LaTEX can only draw lines with certain slopes, TEXcad always takes the line with a proper slope next to the current slope.

3.3.2 EMLines

If you are using one of the emTEX drivers then it is possible to draw lines in any direction, not just these allowed by LaTEX. To do this you must first select the EMLines option in the Options submenu (see section 6.4). You must also add the style file emlines.sty to your document as, for instance,

\documentstyle[emlines,11pt]{article}. emlines.sty contains user macros for drawing lines:

\emline#1#2#3#4#5#6 places point #3 at the coordinates (#1,#2) and point #6 at coordinates (#4,#5) and draws a line between these points. TEXcad generates a call to this macro for each EMLine it draws.

Note: if there is more than one picture to a page then the \emline macro must be preceded by a call to \newpic{digit}. This call causes digit to be placed before the number designating all the #3 and #6 points so that they remain unambiguous.

With V1.3j or later of the emTEX drivers, this redefining of \emline is unnecessary, you must then use emlines2.sty instead. The \emline macro defined there ignores the pointnumbers, so that old pictures can still be used.

3.4 Circles and filled circles

Circles are drawn by selecting Circle and filled circles by selecting Filled Circle. In both cases the centre of the circle is marked first followed by any point on its periphery. As circles in LaTEX can only take certain sizes the “rubber circle” will only follow the cursor when it falls on a permitted radius.

3.5 Ellipses

The items Oval and Half/Quarter Oval allow you to draw ellipses or parts of ellipses. To draw an ellipse you must first mark the bottom lefthand corner of the rectangle enclosing it and then the top righthand corner. For Half/Quarter Oval you will first have to select the part you want from a further submenu.

3.6 Bezier-Curves

The last item, Bezier, allows you to draw bezier-curves by specifying three points. When processing the document with LaTEX, You have to add the style file bezier.sty as, for instance,

\documentstyle[bezier,emlines,11pt]{article}. bezier.sty contains the appropriate macro for

drawing the curves, the file should be available on every LaTEX installation.

4 Editing pictures

The second item on the main menu is Edit, selecting this gives a submenu to choose from:

                                      EDIT

Text-Edit Copy Move Delete Delete all Save Macro Load Macro

Editing allows objects to be copied, moved or deleted. It also allows the whole picture to be deleted, or parts of the picture to be saved or loaded. All these operations, except Delete all, require you to pick the objects to be edited.

4.1 Picking objects

When an operation other than Delete all or Load Macro is selected the following submenu is displayed:

                              COPY/MOVE/DELETE/SAVE MACRO

Pick objects Pick area Unpick

By selecting Pick objects you can select single objects in the picture. This is done by selecting a point depending on the object as given in the following table:

Object  Point
Text  a point near T
Boxes, ellipses a point inside the box or ellipse
Lines, arrows a point approximately on the line
Circles  a point approximately on the periphery
Bezier  a point approximately on the start- or endpoint of the curve

The selected object is marked and you are asked to confirm your choice: if you answer “Yes”, the object is picked and you can then go on to pick other objects or return to the Pick menu by pressing the righthand mouse button. If you answer “No” TEXcad will look to see if there is another object nearby, if so then you will be asked to confirm this choice and so on.

If no object can be found, TEXcad will issue an appropriate message.

If you select Pick area then you will be asked to mark the lower left and upper right corners of a window: all objects which fall completely within this window will be picked and will be so marked.

Selecting Unpick undoes all pick operations.

Important note: if you leave the pick menu with the righthand mouse button then the pending operation will not be terminated, but it still can be terminated thereafter.

4.2 Editing text

By selecting the menu item Text-Edit you can either alter text which already exists in the picture or you can put text into boxes now empty.

First you will be asked to pick either a text marker T or a box. Text input or editing is done in the same way as in section 3.1.

4.3 Copying and moving objects

By selecting Copy or Move from the Edit menu you can copy or move a part of the picture.

After you have picked the objects taking part in the operation, you will be asked to mark the start and end of a vector in the desired movement direction. With Move the picked objects will be placed in the given position and the originals deleted. Copy does the same but does not delete the originals.

4.4 Deletion

Delete is used to remove single objects from the picture, after picking them, TEXcad will ask for confirmation.

Delete all clears all the objects from the picture, you are asked to confirm this too.

4.5 Macros

TEXcad allows you to save parts of the picture and load them into other pictures. These macros are saved as a LaTEX picture environment in the same way as complete pictures – they can therefore be included in documents without further processing. Also can whole picture be read in as macros.

To save a macro, select Save Macro from the EDIT menu. After you have picked the objects to be saved, TEXcad asks for the name of a file in which to write them. When this has been given the macro will be saved. If the name is not a path name then the file will be put in the default directory as shown in Options/Directories. If the filename is given without an extension then the default extension will be used (see section 6).

To load a macro, select Load Macro from the EDIT menu. TEXcad will ask for the name of the file and will then load the picture in it. If no path or extension is given then the file name will be extended with the defaults as for Save Macro. Section 7 contains more information on inputting the file name and the format of the file’s contents. If the file does not exist an appropriate message will be issued. The objects loaded are marked and can be moved to the desired place in the same way as described in section 4.3.

5 Utilities

As well as drawing and editing pictures, TEXcad has a number of utilities to help in the task.

5.1 Redraw

Redraw in the main menu redraws the picture which may be necessary after drawing circles during which some part of the picture may be overwritten.

5.2 The snap function

When snap is enabled, the cursor no longer moves smoothly but jumps from point to point on a grid. Setting snap is done by selecting Snap from the main menu:

                             SNAP (On, < asp >)/SNAP (Off)

On Off Aspect

The function is turned on or off by selecting On or Off. Selecting Aspect allows you to alter the grid spacing. The status of the snap function is displayed in the snap menu heading.

While entering points in the drawing window, the snap function can be toggled on and off with P.

5.3 Zoom

Zoom increases the displayed size of the picture. On selecting the Zoom item from the menu the following submenu appears:

                                    ZOOM (< fac >)

Factor Area All Last

Selecting Factor from this submenu allows you to choose the factor by which the picture will be magnified. On selecting Area you will be asked to mark a window and the magnification will be chosen so that the picture fills this window. The current zoom factor is shown in the zoom menu heading. All sets zoom to one and puts point (0,0) in the bottom lefthand corner of the drawing window, all the picture is displayed. When the magnification is changed, TEXcad saves the last zoom factor and the position of the part of the picture displayed, by selecting Last you can return to the previous zoom/position, switching between two different views.

5.4 Moving the window

Selecting Window from the main menu displays this submenu:

WINDOW Lower left Upper right Origin

By selecting Lower left and marking a point, you can move the displayed image so that the point chosen appears in the lower lefthand corner of the drawing window. In the same way, selecting Upper right puts the selected point at the upper righthand corner of the window. Selecting Origin puts the origin (0,0) in the lower lefthand corner of the window.

While marking points in the window, the E, X, S and D keys can be used to move the display. E and X move the window up and down by a quarter of the window height, S and D move it to the left and right by a quarter of the width. This function is enabled only when a single point is being input, for instance, the starting point of a line, but not when a rubber band is displayed as when entering the end point of a line.

6 Options

Selecting Options in the main menu displays the following submenu:

                                      OPTIONS

Directories

Snap Zoom Unitlength EMLines Save Options Load Options

This submenu allows you to set the options described in section 2.3.

6.1 Directories

Selecting this item displays another submenu:

                                    DIRECTORIES

Picture-Directory Picture-suffix Macro-Directory Macro-Suffix BGI-Directory

Picture-Directory is the directory in which pictures or searched for and stored when the picture file name is not a path name, Picture-Suffix is the default extension for picture file names to be used when none is given. Macro-Directory and Macro-Suffix perform the same functions for macro files.

Selecting BGI-Directory allows the directory containing the Borland graphics driver files to be set.

6.2 Snap, Zoom

These two items have the same function here as in the main menu: they are repeated here to enable the zoom factor and the snap function values to be set as they are values which can be saved.

6.3 Dimensional units

The option Unitlength allows the LaTEX picture parameter \unitlength to be set. It is the unit in terms of which all measurements in the picture environment are expressed.

6.4 EMLines

The option EMLines displays the submenu:

                         EMLINES (Off)/EMLINES (On, < width >)

On Off Linewidth

To be able to draw lines at any angle, On of these option must be selected. Selecting Off will return to the normal LaTEX situation. DVI files with EMLines can only be printed using the emTEX drivers. The option Linewidth allows the thickness of EMLines or of vertical and horizontal lines in the normal LaTEX mode. The current value is displayed in the header of the EMLines menu.

6.5 Loading and saving options

The menu item Save Options saves the current option settings, Load Options sets the options from a parameter file. Defaultname for the parameter file is homedir\texcad.opt, where homedir is the directory texcad.exe was loaded from. When loading a picture path\name.pic, TEXcad loads the parameterfile path\name.opt, if it is present.

7 Loading and saving pictures

The last two items in the main menu allow complete pictures to be loaded and saved. On selecting Save you will be asked for a file name in which the picture is to be saved, Load asks for a file name from which the picture is to be loaded. In both cases, the file name will have the default picture path and extension added to it if either path or extension is not given. If Load is selected when the current picture has not been saved, you will be asked to confirm the operation: if the answer is “Yes” then the current picture will be deleted and replaced with the loaded one.

If the file name given in response to the Load query contains wildcard characters (*, ?), a list of matching filenames will be displayed in the menu window. You can move through this list by selectiong the first or last item (row of up/down arrows), to select the highlighted file press either the lefthand mouse button or the ENTER key – to end the operation without loading a file, press the righthand mouse button or the ESC key.

7.1 Picture file format

Picture files generated by TEXcad have the following format (if no objects have been drawn or loaded before, no file is created):

     \unitlength=length1
     \special{em:linewidth length2}
     \linethickness{length2}
     \begin{picture}(x,y)

\put(x,y){picture-object}

⋮

     \end{picture}

Where EMLines or Bezier-curves are used, the \put command is replaced by the following macro calls:

\emline{x1}{y1}{point-no}{x2}{y2}{point-no}

     \bezier{num}(x1,y1)(x2,y2)(x3,y3)

If you want to modify a picture file or create one from scratch, you should note the following. Everything before and after the picture environment is ignored. Lines within the picture environment which do not contain a \put, \emline or \bezier command are read in and stored unchanged; you can therefore insert any (appropriate) declarations into a completed picture.

The parser which reads pictures when loading is a little fragile: if it finds an error, TEXcad ends after giving you the chance to save the current picture (this will probably be improved in a later version). However, picture files generated by TEXcad will certainly be read in without error.

8 Ending the run

The final option in the main menu ends the run. If the current picture has been modified since it was last saved, you will be asked if you really want to leave TEXcad.

A The file gh.bat

This batch file is intended to save you from typing the same DOS-commands over and over again, when developing a LaTEX-document. gh.bat is called via

                                        gh <file>

where file is the root file of your document and must be entered without .tex. The following menu will be shown:

                             gh.bat 2.0:   Rootfile=file.tex

                               1 = Edit     4 = TeXcad
                               2 = LaTeX    5 = Drucken
                               3 = Preview  0 = Quit

gh.bat calls the appropriate batch files provided with emTEX upon entering one of the displayed numbers. Of course, you may edit the file to adapt it to your own needs.

B The file gh.tex

This file is intended for testing TEXcad picture files. If you run it through LaTEX, the following prompt is displayed:

Enter name of (picture)file:

You can now enter the name of a file, which will then be processed by LaTEX.

C Licensing conditions

TEXcad is shareware, and the conditions of use are:

• TEXcad can be obtained either direct from the author or as included in the emTEX package from Eberhard Mattes. To get it from the author, please send a 5\frac{1}{4} inch double density diskette with return postage to the address below – please send the disk in a container which can be used for the return journey.

• TEXcad can, and should, be copied and passed on but only under the condition that it is completely unchanged (both the files and this manual) and no charge is made. May whoever breaks these conditions be plagued by a bad conscience for the rest of their life!

• Many people have been pleased with TEXcad, and a number of suggestions for improvement have reached me. Anybody who finds the program useful or who would like to learn of later versions should send or transfer 30.– Deutschmarks to me. They will then be informed of improvements and will be sent the latest version. Planned modifications/improvements/extensions are:

– Drawing circles of any diameter – only possible with modifications to the emTEX DVI drivers, perhaps too difficult as drawing circles is not easy.

– Modifying objects, for instance, changing the diameter of a circle.

– Different line thicknesses in the same picture.

– Improvements to the picture environment parser

The author of TEXcad is:

Georg Horn Burgweg 1a 5400 Koblenz-Moselweiß

Bank: Volksbank Hamm/Sieg, BLZ. 573 915 00, Account no. 33.1130.05